# Beyond Content Relevance: Evaluating Instruction Following in Retrieval Models

### v1

@@ Abstract
Instruction-following capabilities in large language models (LLMs) have significantly progressed, enabling more complex user interactions through detailed prompts. However, retrieval systems have not matched these advances, most of them still relies on traditional lexical and semantic matching techniques that fail to fully capture user intent. Recent efforts have introduced instruction-aware retrieval models, but these primarily focus on intrinsic content relevance, which neglects the importance of customized preferences for broader document-level attributes. This study evaluates the instruction-following capabilities of various retrieval models beyond content relevance, including LLM-based dense retrieval and reranking models. We develop InfoSearch, a novel retrieval evaluation benchmark spanning six document-level attributes: Audience, Keyword, Format, Language, Length, and Source, and introduce novel metrics – Strict Instruction Compliance Ratio (SICR) and Weighted Instruction Sensitivity Evaluation (WISE) to accurately assess the models’ responsiveness to instructions. Our findings reveal that while reranking models generally surpass retrieval models in instruction following, they still face challenges in handling certain attributes. Moreover, although instruction fine-tuning and increased model size lead to better performance, most models fall short of achieving comprehensive instruction compliance as assessed by our benchmark. [^1] https://github.com/EIT-NLP/InfoSearch
@@ 1 Introduction
The advent of instruction-following in large language models (LLMs) has greatly expanded their generative capabilities (Lou et al., 2023; Brown, 2020), allowing users to express more complex intentions through detailed instructions (Touvron et al., 2023; Black et al., 2022). However, advances in retrieval systems have not kept pace with these improvements in generative models. Many retrieval systems still rely heavily on traditional lexical or semantic matching techniques (Xiong et al., 2020; Xiao et al., 2023; Wang et al., 2022), which are often too limited to fully capture user intent. As a result, while users have grown accustomed to interacting with generative models using intricate instructions (Team et al., 2023; Bai et al., 2023; Achiam et al., 2023), their sophisticated search behaviors in retrieval remains constrained-primarily centered around keyword-based queries followed by manual filtering of results to find the desired information. Several studies have started to explore instruction-aware retrievers that can interact with users as seamlessly as generative models, but they mainly focus on task-level instructions (Asai et al., 2023; Wang et al., 2023; Su et al., 2023; Peng et al., 2024), guiding retrievers with one instruction for each task. While this task-level instruction is essential for adapting a single retrieval model to multiple predefined scenarios, it falls short of meeting users’ customized demands beyond standard tasks Weller et al. (2024b).
@@
Recently, a few works have started shifting from task-level to instance-level instructions, providing tailored instructions for each instance to better reflect customized needs (Weller et al., 2024a; Oh et al., 2024). These efforts incorporate instructions that specify which content to include or exclude, thus clarifying user intent. While they greatly enrich the diversity of instructions, their primary focus remains on content relevance. Typically when searching for certain documents, users care about two types of aspects: the informational content they provide and the manner in which it is presented (Taylor, 1962; Mizzaro, 1998), which involves document-level attributes such as length, language, and format. A sole focus on content relevance neglects the importance of customized preferences for broader document-level attributes.
@@
We believe that a truly instruction-aware retrieval system must go beyond content relevance to accommodate a range of diverse, user-defined document attributes. To further research in this direction, we propose InfoSearch, a novel benchmark designed to evaluate IR models based on their ability to follow customized instructions across six structured dimensions: Audience, Keyword, Format, Language, Length, and Source. These dimensions reflect a wide range of document-level features that are crucial to fulfilling user requirements beyond content relevance. In addition, we include both instructed and reversely instructed modes to evaluate the model’s comprehension of instructions presented in affirmative and negative formats. Each instruction is carefully crafted and manually validated to ensure naturalness and representativeness of complex real-world scenarios.
@@
Beyond the comprehensiveness of datasets, well-defined evaluation metrics are crucial for gaining a thorough understanding of the instruction-following capabilities of retrieval models. While traditional IR metrics like nDCG and MRR are primarily effective for assessing content relevance in ad-hoc retrieval tasks Weller et al. (2024a), we propose new metrics specifically designed to measure the depth of instruction-following capabilities in retrieval models. By structuring the evaluation across these separate dimensions and modes, we offer a detailed analysis of how well models follow instructions on each condition, providing clearer insights into their strengths and limitations. Overall, instruction-specific fine-tuning and increased model parameters enhance the model’s instruction-following capabilities and re-ranking models demonstrating greater proficiency than retrieval models in instruction-following. Nevertheless, both methods still exhibit substantial potential for improvement in satisfying our benchmark.
@@
  • We introduce two novel metrics – Strict Instruction Compliance Ratio (SICR) and Weighted Instruction Sensitivity evaluation (WISE) metric, to evaluate instruction adherence both strictly and leniently.
  • We evaluate 15 retrieval models, including 1 sparse retrieval method, 8 dense models, and 6 LLM-based reranking models. This extensive evaluation provides a comprehensive comparison of different methodologies, offering valuable insights into their instruction-following capabilities.
@@ 2 InfoSearch: Construction and Evaluation
We construct a benchmark, Instruction-Following Search (InfoSearch), to evaluating the search models’ ability to follow instructions. The benchmark is composed of instructional query/doc data across six dimensions and two novel metrics to measure models’ responsiveness to instructions. Section §2.1 details the dimension settings and retrieval modes under the InfoSerach framework, Section §2.2 explains the dataset construction process, and Section §2.3 describes the design logic behind our proposed metrics.
@@ 2.1 Dataset Framework
In real-world scenarios, users exhibit a wide range of complex and diverse search intentions. The use of tailored instructions can link the specific search query content to the requirements and preferences of the user. Instructions typically contain detailed information or document-level characteristics that align with user needs, aiming to enhance the precision and relevance of search results. As shown in Figure 2.1, we conducted an extensive analysis of real users and their underlying intentions to identify six factors (dimensions) influencing search behaviors: user background (Audience), specific search terms or topics (Keyword), preferred format for information presentation (Format), required response length (Length), language requirement (Language), and information origin (Source). To enhance the diversity of instructions across these six dimensions, we established multiple conditional branches within each dimension, allowing instructions to dynamically adapt and expand based on different conditions. Moreover, even within the same dimension and conditions, we create varied instructions using diverse wording and expressions. This approach not only enriches the dataset but also strengthens the robustness and reliability of the evaluation framework by simulating a broad range of potential user inputs.
Drawing inspiration from Zhang et al. (2024), we incorporate semantic negation into the dataset by reversing the meaning of instructions across each dimension. This approach allows each query to be associated with multiple instructions, covering various conditions and offering both positive and negative semantic contexts. This ensures that the model is exposed to three distinct retrieval modes:
@@ 2.2 Construction Process
The primary objective of developing InfoSearch is to establish a linkage between queries and diverse instructions, ensuring that the corresponding target documents are aligned to each specific instruction. As shown in Figure 2, to achieve this, we specifically gather Question-Answer (Q-A) pairs corresponding to each dimension and expand the set of target documents by incorporating relevant content retrieved from web pages. The specific steps for constructing the framework are as follows:
@@
Dimension # $Q$ Avg $|Q|$ $I$ Avg $|I|$ $R$ Avg $|R|$ $D$
Audience 100 9.02 210 20.46 210 15.91 840
Keyword 100 6.30 288 17.90 288 18.92 1152
Format 100 9.16 300 16.65 300 19.31 1200
Language 100 8.75 200 14.09 200 15.74 800
Length 100 8.52 300 15.94 300 16.26 1200
Source 100 7.38 300 18.19 300 15.58 1200
Total 600 1598 1598 6392
@@ 1. Step 1:
    Condition Determination: The query for each dimension can be diversified and expanded through various conditions, allowing the same query to correspond to different relevant documents under different conditions. Except for the Keyword dimension, the other five dimensions have fixed conditions. The condition of the Keyword dimension is the keyword in the document that is relevant to the query. Therefore, the condition of the Keyword dimension needs to be determined after filtering out the Query-Document (Q-D) pairs. To achieve this, both the query and its corresponding document were input into GPT-4, generating a unique condition for each Q-D pair. [^2] GPT-4 occasionally selected irrelevant words, such as “and” or “what”, that did not align with the user scenario when generating keyword conditions. To address this, we meticulously crafted prompt templates for condition extraction, ensuring that the conditions were both unique and representative of each document, accurately reflecting the document’s core theme.
@@ 2. Step 2:
    Data Collection: To ensure the naturalness of the queries and reduce production costs, we consciously integrate conditions when filtering Q-A pairs from existing datasets or web pages. For instance, in the Format dimension, due to the lack of available multi-format Q&A datasets, we selectively extract Q-D pairs from StackOverflow posts. For posts containing code and detailed official documentation responses, we use their titles as queries and the complete responses as documents under the [StackOverflow] condition. The pure code snippets within the answers and references to official documentation are separately extracted and used as documents under the [Code Snippet] and [Manual] conditions, respectively.
@@ 3. Step 3:
    Instruction Generation: The generation of accurate, concise, and natural instructions is crucial. When searching, users tend to express their intentions using simple, naural language, so the generated instructions must remain brief and clear, closely mirroring conversational style. To achieve this, we employed words such as “smooth”, “natural”, and “realistic” in the prompts to guide GPT-4 in crafting instructions that emphasize not only semantic accuracy but also the emulation of authentic user expressions. Furthermore, a two-sentence structure for the instructions, first rephrasing the query and then appending specific constraints. This structure effectively separates the core query from the conditions, enhancing the diversity of generated instructions.  [^3] For example, “What are the most effective exercises for losing weight? Please find discussions from forum posts only.” This two-sentence structure ensures logical clarity and semantic coherence.
@@ 4. Step 4:
    Document Rewriting: When the query and relevant documents fail to meet the instruction requirements, we use GPT-4 to rewrite the existing documents to generate relevant documents. In this process, we experimented with directly generating condition-satisfying documents from the query, but these often exhibited redundant expressions and inconsistent formatting. Therefore, we adjusted the existing documents, ensuring they meet the instruction requirements while preserving naturalness and authenticity in the language.
@@ 5. Step 5:
    Instruction Reversal: In real retrieval, we observed that results already ranked highly tend to remain at the top even after instructions are applied. This makes it difficult to determine whether the improvement in ranking is due to the model’s understanding of the instructions or simply a result of detailed keyword and semantic matching. To address this, we validate the model’s comprehension of the instructions by reversing the semantic meaning of the instructions. For example, “Please answer in Chinese” is reversed to “Please do not answer in Chinese.”
@@ 6. Step 6:
    Hard Negative Generations: While positive documents for the same query under varying conditions may act as negative examples for one another (instruction negatives), we still need to prevent the model from relying solely on prominent features for simple retrieval, thereby neglecting the subtle relationships between the query and the documents. To address this, we use GPT-4 to generate documents that appear to be related to the query topic on the surface but cannot actually answer the query, serving as hard negative documents (query negatives).
@@ 7. Step 7:
    Manual Review: We filtered out anomalous documents from the outputs of 12 retrieval models, selecting those that failed to rank within the top 50 in six or more models or had a relevance score below 0.5 for the query, followed by manual screening. This process aimed to eliminate mislabeled Q-D pairs selected from other datasets or documents inaccurately retrieved through manual search. For these mismatched documents, we proceed with manual replacement.
By applying the data construction process described above, the InfoSearch benchmark comprises 600 core queries, 1,598 instructed queries, 1,598 reversely instructed queries, and 6,392 documents, as summarized in Table 1. Detailed examples of each dimension are provided in Appendix A.
@@ 2.3 Evaluation Metrics
$min\mbox{-nDCG}(q_{i})=\min_{j\in(1,2,...,m_{i})}s_{ij}$ (1)
$Robustness@k=\frac{1}{n}\sum_{i=1}^{n}min\mbox{-nDCG}(q_{i})$ (2)
[…]
Notation: To address these shortcomings, we define two metrics to quantify the model’s responsiveness to instructions: Strict Instruction Compliance Ratio (SICR) and Weighted Instruction Sensitivity Evaluation (WISE) metric, guided by the practical demands of real-world retrieval systems. Assuming that in the original mode, the core query $q$ has $n$ positive documents, denoted as ${\color[rgb]{0,0,0}\textbf{P}}=\{P_{1},P_{2},\ldots,P_{n}\}$. When it comes to the instructed mode where the core query is accompanied by a more specific instruction, there is only one unique gold document $\textbf{P}^{i}$out of P. The remaining documents become negative. When it comes to the Reversely instructed mode, as the instruction is semantically reversed, $\textbf{P}^{i}$  becomes negative. Let $R_{ori}$, $R_{ins}$ and $R_{rev}$ represent the rankings of the gold document $\textbf{P}^{i}$ in original mode, instructed mode, and reversely instructed mode, respectively. Similarly, let $S_{ori}$, $S_{ins}$ and $S_{rev}$ denote the relevance score between the gold document $\textbf{P}^{i}$ and the query in original mode, instructed mode, and reversely instructed mode, respectively.
@@
When defining the reward component, the model is expected to comprehend and execute the instructions, effectively optimizing the rankings of the top K results accordingly. This implies that significant ranking changes within the top K results should be given greater weight to address defects ③ and ④, as these changes are more likely to be noticed and utilized by users. It is essential to consider both the absolute ranking $R_{ins}$ and the relative ranking $(R_{ori}-R_{ins})$. To achieve this, we introduced the $\frac{1}{\sqrt{R_{ins}}}$ term, generously rewarding smaller $R_{ins}$ values. Simultaneously, through the $(1-\frac{R_{ori}-R_{ins}}{K})$ term, we grant higher rewards to results that demonstrate substantial ranking improvements. Additionally, a uniform value of 0.01 is assigned to results beyond the Top K. Additionally, more extreme cases are considered (defects ⑤): if a core query contains $N$ positive documents in the original mode and meets the conditions $R_{ori}\leq N$ and $R_{ins}=1$, it will be granted a reward of 1. This is based on the premise that, for an ideal retriever, the $N$ positive documents would likely rank at the top in the original mode. Accordingly, results that rank higher and exhibit more significant changes should be assigned greater weight. The reward formula is defined as:
$f_{reward}=\begin{cases}1,&\text{if}~{}R_{ori}\leq N~{}\text{ and }~{}R_{ins}=1,\\ (1-\frac{R_{ori}-R_{ins}}{K})\cdot\frac{1}{\sqrt{R_{ins}}},\quad\quad&\text{if}~{}R_{ori}\leq K,\\ 0.01,&\text{otherwise},\end{cases}$ (7)
where $K=20$ signifies that our primary focus is on the top 20 retrieval results.
@@ 3.1 Experimental Setup
Based on Mistral-7B-Instruct-v0.2 Jiang et al. (2023), we conduct a specialized experiment to evaluate its zero-shot performance in three retrieval settings: dense retrieval, point-wise reranking, and list-wise reranking. As a highly capable instruction-tuned model, Mistral is expected to demonstrate instruction-following abilities in retrieval tasks without fine-tuning. This experiment explores Mistral’s potential as a zero-shot retrieval model, assessing whether it can naturally generate strong embeddings or act as an effective reranker to identify instruction-relevant documents from the list of candidates. For the mode of dense retrieval, we use mean pooling to obtain the sentence level representation.
@@
List-wise Reranking
Mistral-ins-v0.2 74.5 64.4 61.6 74.5 40.5 42.2 7.2 8.1 22.0 5.7 4.8 7.2

### v2

@@ Abstract
Instruction-following capabilities in LLMs have progressed significantly, enabling more complex user interactions through detailed prompts. However, retrieval systems have not matched these advances, most of them still relies on traditional lexical and semantic matching techniques that fail to fully capture user intent. Recent efforts have introduced instruction-aware retrieval models, but these primarily focus on intrinsic content relevance, which neglects the importance of customized preferences for broader document-level attributes. This study evaluates the instruction-following capabilities of various retrieval models beyond content relevance, including LLM-based dense retrieval and reranking models. We develop InfoSearch, a novel retrieval evaluation benchmark spanning six document-level attributes: Audience, Keyword, Format, Language, Length, and Source, and introduce novel metrics – Strict Instruction Compliance Ratio (SICR) and Weighted Instruction Sensitivity Evaluation (WISE) to accurately assess the models’ responsiveness to instructions. Our findings indicate that although fine-tuning models on instruction-aware retrieval datasets and increasing model size enhance performance, most models still fall short of instruction compliance. [^1] We release our dataset and code on https://github.com/EIT-NLP/InfoSearch
@@ 1 Introduction
The advent of instruction-following in large language models (LLMs) has greatly expanded their generative capabilities (Brown, 2020; Lou et al., 2023), allowing users to express more complex intentions through detailed instructions (Black et al., 2022; Touvron et al., 2023). However, retrieval systems have not kept pace with these advancements, continuing to rely on traditional lexical or semantic matching techniques (Xiong et al., 2020; Wang et al., 2022; Xiao et al., 2023). As a result, while users have grown accustomed to interacting with generative models using intricate instructions (Team et al., 2023; Bai et al., 2023; Achiam et al., 2023), their retrieval behavior remains limited to keyword-based queries followed by manual filtering of results to find the desired information. Several studies have started to explore instruction-aware retrievers that can interact with users as seamlessly as generative models, but these primarily focus on task-level instructions (Asai et al., 2023; Wang et al., 2023; Peng et al., 2024), guiding retrievers with one instruction for each task. While this task-level instruction is essential for adapting a single retrieval model to multiple predefined scenarios, it falls short of meeting users’ customized demands beyond standard tasks (Weller et al., 2024b).
@@
Recent works have shifted from task-level to instance-level instructions, providing tailored instructions for each instance to better align with customized needs (Weller et al., 2024a; Oh et al., 2024). These approaches specify instructions that which content to include or exclude, thus clarifying user intent. While they greatly enrich the diversity of instructions, their primary focus remains on content relevance. When searching for certain documents, users typically care about two aspects: the informational content and its presentation (Taylor, 1962; Mizzaro, 1998), including document-level attributes such as length, language, and format. A sole focus on content relevance neglects the importance of customized preferences for broader document-level attributes.
@@
We believe that a truly instruction-aware retrieval system must go beyond content relevance to accommodate a variety of user-defined document attributes. To further research in this direction, we propose InfoSearch, a novel benchmark designed to evaluate IR models based on their ability to follow customized instructions across six structured dimensions: Audience, Keyword, Format, Language, Length, and Source. These dimensions encompass key document-level features that address user needs beyond content relevance. Additionally, we include both instructed and reverse-instructed modes to assess the model’s ability to comprehend instructions in both affirmative and negative formats. Each instruction is carefully crafted and manually validated to ensure naturalness and representativeness of complex real-world scenarios.
@@
Beyond the comprehensiveness of datasets, well-defined evaluation metrics are essential to thoroughly assess the instruction-following capabilities of retrieval models. While traditional IR metrics like nDCG and MRR are primarily effective for assessing content relevance in ad-hoc retrieval tasks (Weller et al., 2024a), we propose new metrics specifically designed to measure the depth of instruction-following capabilities in retrieval models. By structuring the evaluation across these separate dimensions and modes, we offer a detailed analysis of how well models follow instructions on each condition, providing clearer insights into their strengths and limitations. Overall, fine-tuning on instruction-aware retrieval datasets and increasing model parameters improve instruction-following capabilities, with re-ranking models outperforming retrieval models in this aspect. However, both approaches still show considerable room for improvement in meeting the standards set by our benchmark.
@@
  • We introduce two novel metrics – Strict Instruction Compliance Ratio (SICR) and Weighted Instruction Sensitivity Evaluation (WISE) – that provide a more nuanced and accurate assessment of retrieval models’ instruction adherence compared to traditional IR metrics.
  • We evaluate 15 retrieval models, encompassing 1 sparse retrieval model, 8 bi-encoder-based dense models, and 6 LLM-based reranking models. This thorough evaluation enables a comprehensive comparison across diverse methodologies, delivering valuable insights into their instruction-following effectiveness.
@@ 2 InfoSearch: Construction and Evaluation
We construct a benchmark, Instruction-Following Search (InfoSearch), to evaluating the search models’ ability to follow instructions. InfoSearch is composed of query-doc pairs across six dimensions and two novel metrics to measure models’ responsiveness to instructions. In this section, §2.1 details the dimension settings and retrieval modes under the InfoSerach framework, §2.2 explains the dataset construction process, and §2.3 describes the design logic behind our proposed metrics.
@@ 2.1 Dataset Framework
In real-world scenarios, users exhibit a wide range of complex and diverse search intentions. The use of tailored instructions can link the specific search query content to the requirements and preferences of the user. Instructions typically contain detailed information or document-level characteristics that align with user needs, aiming to enhance the precision and relevance of search results. As shown in Figure 1, we conducted an extensive analysis of real users and their underlying intentions to identify six factors (dimensions) influencing search behaviors: user background (Audience), specific search terms or topics (Keyword), preferred format for information presentation (Format), required response length (Length), language requirement (Language), and information origin (Source). To enhance the diversity of instructions across these six dimensions, we established multiple conditional branches within each dimension, allowing instructions to dynamically adapt and expand based on different conditions. Moreover, even within the same dimension and conditions, we create varied instructions using diverse wording and expressions. This approach not only enriches the dataset but also strengthens the robustness and reliability of the evaluation framework by simulating a broad range of potential user inputs.
Drawing inspiration from (Zhang et al., 2024), we incorporate semantic negation into the dataset by reversing the meaning of instructions across each dimension. This approach allows each query to be associated with multiple instructions, covering various conditions and offering both positive and negative semantic contexts. This ensures that the model is exposed to three distinct retrieval modes:
@@ 2.2 Construction Process
The primary objective of developing InfoSearch is to bridge queries with diverse instructions, ensuring precise alignment between instructions and their corresponding target documents. We achieve this by collecting Question-Answer (Q-A) pairs for each dimension and expanding the target document pool through web-retrieved content. Figure 2 outlines the 7-step construction process. Data sources, methodological details (e.g., GPT-4 prompts), implementation challenges and dataset statistics are provided in Appendix A.
@@ 1. Step 1:
    Condition Determination: Queries are diversified via multiple conditions, enabling a single query to yield distinct relevant documents depending on contextual requirements.
@@ 2. Step 2:
    Data Collection: To ensure query naturalness and minimize generation costs, we consciously integrate conditions when filtering Q-A pairs from existing datasets or web pages.
@@ 3. Step 3:
    Instruction Generation: Requires producing precise, concise, and natural instructions that reflect natural conversational patterns, aligning with users’ tendency to express intents.
@@ 4. Step 4:
    Document Rewriting: When queries or documents inadequately address instruction requirements, GPT-4 refines existing content to produce contextually appropriate documents.
@@ 5. Step 5:
    Instruction Reversal: To verify whether ranking improvements stem from instruction comprehension, instruction semantics are systematically reversed, testing model robustness against persistent high-ranking results.
@@ 6. Step 6:
    Hard Negative Generations: Adversarial examples are added to resist the model’s tendency to depend on superficial document features rather than query-document relationships.
@@ 7. Step 7:
    Manual Review: Anomalous outputs were excluded, prioritizing documents that consistently underperformed or had low relevance scores, followed by expert verification.
By applying the data construction process described above, the InfoSearch benchmark comprises 600 core queries, 1,598 instructed queries, 1,598 reversely instructed queries, and 6,392 documents.
@@ 2.3 Evaluation Metrics
We define two metrics to quantify the model’s responsiveness to instructions: Strict Instruction Compliance Ratio (SICR) and Weighted Instruction Sensitivity Evaluation (WISE) metric. Assuming that in the original mode, the core query $q$ has $n$ positive documents, denoted as ${\color[rgb]{0,0,0}\textbf{P}}=\{P_{1},P_{2},\ldots,P_{n}\}$. When it comes to the instructed mode where the core query is designates a single gold document $\textbf{P}^{i}$out of P, demoting others to negatives. When In the reversely instructed mode, $\textbf{P}^{i}$  becomes negative. Let $R_{ori}$, $R_{ins}$ and $R_{rev}$ denote $\textbf{P}^{i}$ ’s rankings and $S_{ori}$, $S_{ins}$ and $S_{rev}$ its relevance scores across original, instructed, and reversed modes.
@@
When defining the reward component, the model is expected to comprehend and execute the instructions, effectively optimizing the rankings of the top K results accordingly. This implies that significant ranking changes within the top K results should be given greater weight to address defects ③ and ④, as these changes are more likely to be noticed and utilized by users. It is essential to consider both the absolute ranking $R_{ins}$ and the relative ranking $(R_{ori}-R_{ins})$. To achieve this, we introduced the $\frac{1}{\sqrt{R_{ins}}}$ term, generously rewarding smaller $R_{ins}$ values. Simultaneously, through the $(1-\frac{R_{ori}-R_{ins}}{K})$ term, we grant higher rewards to results that demonstrate substantial ranking improvements. Additionally, a uniform value of 0.01 is assigned to results beyond the Top K. More extreme cases are considered (defects ⑤): if a core query contains $N$ positive documents in the original mode and meets the conditions $R_{ori}\leq N$ and $R_{ins}=1$, it will be granted a reward of 1. This is based on the premise that, for an ideal retriever, the $N$ positive documents would likely rank at the top in the original mode. Accordingly, results that rank higher and exhibit more significant changes should be assigned greater weight. The reward formula is defined as:
$f_{reward}=\begin{cases}1,&\text{if}~{}R_{ori}\leq N~{}\text{ and }~{}R_{ins}=1,\\ (1-\frac{\sqrt{R_{ori}-R_{ins}}}{K})\cdot\frac{1}{\sqrt{R_{ins}}},\quad\quad&\text{if}~{}R_{ori}\leq K,\\ 0.01,&\text{otherwise},\end{cases}$ (4)
where $K=20$ signifies that our primary focus is on the top 20 retrieval results. Some of the reward values are visualized in the Figure 5.
@@ 3.1 Experimental Setup
Based on Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), we conduct a specialized experiment to evaluate its zero-shot performance in three retrieval settings: dense retrieval, point-wise reranking, and list-wise reranking. As a highly capable instruction-tuned model, Mistral is expected to demonstrate instruction-following abilities in retrieval tasks without fine-tuning. This experiment explores Mistral’s potential as a zero-shot retrieval model, assessing whether it can naturally generate strong embeddings or act as an effective reranker to identify instruction-relevant documents from the list of candidates. For the mode of dense retrieval, we use mean pooling to obtain the sentence level representation.
@@
List-wise Reranking (Fine-tuned)
@@
List-wise Reranking (Instructional Zero-shot)
Mistral-ins-v0.2 74.5 64.4 61.6 74.5 40.5 42.2 7.2 8.1 22.0 5.7 4.8 7.2